### v2

# Controllable Video Generation with Provable Disentanglement

Abstract
Controllable video generation remains a significant challenge, despite recent advances in generating high-quality and consistent videos. Most existing methods for controlling video generation treat the video as a whole, neglecting intricate fine-grained spatiotemporal relationships, which limits both control precision and efficiency. In this paper, we propose Controllable Video Generative Adversarial Networks (CoVoGAN) to disentangle the video concepts, thus facilitating efficient and independent control over individual concepts. Specifically, following the minimal change principle, we first disentangle static and dynamic latent variables. We then leverage the sufficient change property to achieve component-wise identifiability of dynamic latent variables, enabling disentangled control of video generation. To establish the theoretical foundation, we provide a rigorous analysis demonstrating the identifiability of our approach. Building on these theoretical insights, we design a Temporal Transition Module to disentangle latent dynamics. To enforce the minimal change principle and sufficient change property, we minimize the dimensionality of latent dynamic variables and impose temporal conditional independence. To validate our approach, we integrate this module as a plug-in for GANs. Extensive qualitative and quantitative experiments on various video generation benchmarks demonstrate that our method significantly improves generation quality and controllability across diverse real-world scenarios.
1 Introduction
Video generation (vondrick2016generating, ; tulyakov2018mocogan, ; wang2022internvideo, ) has become a prominent research focus, driven by its wide-ranging applications in fields such as world simulators (openai2024video, ), autonomous driving (wen2024panacea, ; wang2023drivedreamer, ), and medical imaging (li2024endora, ; cao2024medical, ). In particular, controllable video generation (zhang2025moonshot, ) is essential for advancing more reliable and efficient video generation models. Despite the impressive results achieved by recent commercial or large-scale models such as Kling kling2024 and Wan wan2024wan , precise control over specific aspects of generated video remains a significant challenge, as illustrated in Figure 1. This issue may arise because these models typically represent the video as a unified 4D spatiotemporal block and apply conditioning signals (e.g., text prompts) directly to this global representation. Recent approaches (ho2022video, ; zhou2022magicvideo, ; yang2024cogvideox, ; opensora, ) follow a similar strategy, with differences in modeling frameworks (e.g., diffusion or VAE) and shapes (e.g., vectors or spatiotemporal blocks). However, these formulations often neglect the intricate spatio-temporal structure of videos, thereby limiting the ability to disentangle and control fine-grained factors, such as head movements and eye blinking in facial video generation.
Figure 1: Videos are generated using Kling and Wan with the prompt: "while this person was speaking, the head gradually shifted from the middle to the right." The first row shows that essential motion cues are partially omitted, while in the second row the head size changes undesirably.
To address this issue, one intuitive solution is to learn a disentangled representation of the video, within which the internal relationships are often not considered. Some models (hyvarinen2000independent, ; tulyakov2018mocogan, ; yu2022generating, ; skorokhodov2022stylegan, ; wei2024dreamvideo, ) explicitly decompose video generation into two parts: motion and identity, representing dynamic and static information, respectively. This separation allows for more targeted control over each aspect, making it possible to modify the motion independently without affecting the identity. zhang2025moonshot ; shen2023mostganv leverage attention mechanisms to further disentangle different concepts within the video, enhancing the ability to control specific features with greater precision. fei2024dysen ; lin2023videodirectorgpt utilize Large Language Models to find the intricate temporal dynamics within the video and then enrich the scene with reasonable details, enabling a more transparent generative process. These methods are intuitive and effective, yet they lack a solid guarantee of disentanglement, making the control less predictable and potentially leading to unintentional coupling of different aspects of the video.
These limitations of previous approaches motivate us to rethink the paradigm of video generation. Inspired by recent advancements in nonlinear Independent Component Analysis (ICA) (hyvarinen2017nonlinear, ; khemakhem2020variational, ; yao2022temporally, ; hyvarinen2016unsupervised, ) and the successful applications like video understanding (chen2024caring, ), we propose the Controllable Video Generative Adversarial Network (CoVoGAN) with a Temporal Transition Module plugin. Building upon StyleGAN2-ADA Karras2020ada , we distinguish between two types of factors: the dynamic factors that evolve over time, referred to as style dynamics, and the static factors that remain unchanged, which we call content elements. We also distinguish between different components within the dynamics, allowing for more precise control. By leveraging the minimal change principle, we demonstrate their block-wise identifiability (von2021self, ; li2024subspace, ) and find the conditions under which motion and identity can be disentangled, explaining the effectiveness of the previous line of video generation methods that separate motion and identity. In addition, we employ sufficient change property to disentangle different concepts of motion, such as head movement or eye blinking. Specifically, we introduce a flow (rezende2015variational, ) mechanism to ensure that the estimated style dynamics are mutually independent conditioned on the historical information. Furthermore, we prove the component-wise identifiability of the style dynamics and provide a disentanglement guarantee for the motion in the video.
We conduct both quantitative and qualitative experiments on various video generation benchmarks. For quantitative evaluation, we use FVD (unterthiner2019fvd, ) to assess the quality of the generated videos. We also compare SAP kumar2017variational , modularity ridgeway2018learning and MCC to demonstrate the disentanglement capability. For qualitative analysis, we evaluate the degree of disentanglement by manipulating different dimensions of the latent variables across multiple datasets and comparing the resulting video outputs. Experimental results demonstrate that our method significantly outperforms other video generation models of a scale similar to CoVoGAN.
Key insights and contributions of our research include:
  • We propose a Temporal Transition Module to achieve a disentangled representation, which leverages minimal change principle and sufficient change property.
  • We implement the Module in a GAN, i.e., CoVoGAN, to learn the underlying generating process from video data with disentanglement guarantees, enabling more precise and interpretable control.
  • To the best of our knowledge, this is the first work to provide an identifiability theorem in the context of video generation. This helps to clarify previous intuitive yet unproven techniques and suggests potential directions for future exploration.
  • Extensive evaluations across multiple datasets demonstrate the effectiveness of CoVoGAN, achieving superior results in terms of both generative quality and controllability.
2 Problem setup
2.1 Generating process
Consider a video sequence $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{T}\}$ consisting of $T$ consecutive frames. Each frame ${\mathbf{x}}_{t}\in\mathbb{R}^{n_{x}}$ is generated via an arbitrary nonlinear mixing function $g$, which maps a set of latent variables to the observed frame ${\mathbf{x}}_{t}$. The latent variables are decomposed into two distinct parts: ${\mathbf{z}}_{t}^{s}\in\mathbb{R}^{n_{s}}$, capturing the style dynamics that evolve over time, and ${\mathbf{z}}^{c}\in\mathbb{R}^{n_{c}}$, encoding the content variables that remain consistent across all frames of the video. Furthermore, these latent variables are assumed to arise from a stationary, non-parametric, time-delayed causal process.
Figure 2: The generating process. The gray shade of nodes indicates that the variable is observable.
As shown in Figure 2 and Equation 1, the generating process is formulated as:
{$&{\mathbf{x}}_{t}=g({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}),\\ &z_{t,i}^{s}=f^{s}_{i}\left(\mathbf{Pa}(z_{t,i}^{s}),\epsilon_{t,i}^{s}\right),\\ &z_{j}^{c}={f^{c}_{j}\left(\epsilon_{j}^{c}\right)},\end{aligned}\right.\text{with}\left\{\begin{aligned} \epsilon^{s}_{t,i}\sim p^{s}_{\epsilon_{i}},\\ \epsilon^{c}_{j}\sim p^{c}_{\epsilon_{j}},\end{aligned}\right.$ (1)
in which $z_{t,i}^{s},z_{i}^{c}\in\mathbb{R}$ refers to the $i$-th entry of ${\mathbf{z}}_{t}^{s}$ and $j$-th entry of ${\mathbf{z}}^{c}$, respectively. $\mathbf{Pa}(z_{t,i}^{s})$ refers to the time-deleyed parents of $z_{t,i}^{s}$. All noise terms $\epsilon_{t,i}$ and $\epsilon_{j}$ are independently sampled from their respective distributions: $p_{\epsilon_{i}}^{s}$ for the $i$-th entry of the style dynamics, and $p_{\epsilon_{j}}^{c}$ for the $j$-th entry of the content elements. The components of ${\mathbf{z}}_{t}^{s}$ are mutually independent, conditioned on all historical variables $\cup_{i=1}^{n_{s}}\mathbf{Pa}(z_{t,i}^{s})$. The non-parametric causal transition $f_{i}$ enables an arbitrarily nonlinear interaction between the noise term $\epsilon_{t,i}$ and the set of parent variables $\mathbf{Pa}(z_{t,i}^{s})$, allowing for flexible modeling of the style dynamics.
2.2 Identification of the latent causal process
For simplicity, we denote $\textbf{f}^{s}$ as the group of functions $\{f^{s}_{i}\}_{i=1}^{n_{s}}$, and similarly for $\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c}$.
Definition 2.1 (Observational Equivalence).
Let $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{T}\}$ represent the observed video generated by the true generating process specified by $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$, as defined in Equation 1. A learned model $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$ is observationally equivalent to the true process if the model distribution $p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)$ matches the data distribution $p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)$ for all values of $V$.
Illustration. When observational equivalence is achieved, the distribution of videos generated by the model exactly matches that of the ground truth, i.e., the training set. In other words, the model produces video data that is indistinguishable from the actual observed data.
Definition 2.2 (Block-wise Identification of Generating Process).
Let the true generating process be $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$ as specified in Equation 1 and let its estimation be $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$. The generating process is identifiable up to the subspace of style dynamics and content elements, if the observational equivalence ensures that the estimated $(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$ satisfies the condition that there exist bijective mappings from $(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$ to $({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c})$ and from $\hat{{\mathbf{z}}}^{c}$ to ${\mathbf{z}}^{c}$. Formally, there exists invertible functions $h:\mathbb{R}^{n_{s}+n_{c}}\rightarrow\mathbb{R}^{n_{s}+n_{c}}$ and $h_{c}:\mathbb{R}^{n_{c}}\rightarrow\mathbb{R}^{n_{c}}$ such that
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)$ (2)
$\Rightarrow$ $\ \ [{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}]),{\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}^{c}),$
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation.
Illustration. When block-wise identification is achieved, content elements are effectively disentangled. As a result, motion control can be applied independently, allowing manipulation of motion without altering the video’s content. For example, in a video where a camera moves forward, it becomes easy to change the camera’s direction without affecting the scene itself.
Definition 2.3 (Component-wise Identification of Style Dynamics).
On top of Definition 2.2, when $h^{s}$ is a combination of permutation $\pi$ and a component-wise invertible transformation $\mathcal{T}$. Formally,
$\ \ p_{(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})}(V)=p_{(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})}(V)$ (3)
$\Rightarrow$ $\ \ {\mathbf{z}}_{t}^{s}=(\pi\cdot\mathcal{T})(\hat{g}^{-1}_{s}({\mathbf{x}}_{t})).$
Illustration. Achieving component-wise identification ensures that each estimated style variable corresponds exactly to one true style variable, which facilitates the disentangling of motion features. This enables efficient and precise control over video generation. For example, when generating a video of a person’s face, one can control different aspects, such as head movements or eye blinks, by adjusting the corresponding variables.
3 Theoretical analysis
In this section, we discuss the conditions under which the block-wise identification (Definition 2.2) and component-wise identification (Definition 2.3) hold.
3.1 Block-wise identification
Without loss of generality, we first consider the case where $\text{Pa}(z_{t,i}^{s})={\mathbf{z}}_{t-1}^{s}$, meaning that the time-dependent effects are governed by the dynamics of the previous time step.
Definition 3.1.
(Linear Operator) Consider two random variables $a$ and $b$ with support $\mathcal{A}$ and $\mathcal{B}$, the linear operator $L_{b\mid a}$ is defined as a mapping from a density function $p_{a}$ in some function space $\mathcal{F}(\mathcal{A})$ onto the density function $L_{b\mid a}\circ p_{a}$ in some function space $\mathcal{F}(\mathcal{B})$,
$\mathcal{F}(\mathcal{A})\rightarrow\mathcal{F}(\mathcal{B}):\ p_{b}=L_{b\mid a}\circ p_{a}=\int_{\mathcal{A}}p_{b\mid a}(\cdot\mid a)p_{a}(a)da.$
Theorem 3.2 (Block-wise Identifiability).
Consider video observation $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{T}\}$ generated by process $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$ with latent variables denoted as ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{c}$, according to Equation 1, where ${\mathbf{x}}_{t}\in\mathbb{R}^{n_{x}},{\mathbf{z}}_{t}^{s}\in\mathbb{R}^{n_{s}},{\mathbf{z}}^{c}\in\mathbb{R}^{n_{c}}$. If assumptions
  • B1 (Positive Density) the probability density function of latent variables is always positive and bounded;
  • B2 (Minimal Changes) the linear operators $L_{{\mathbf{x}}_{t+1}\mid{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}}$ and $L_{{\mathbf{x}}_{t-1}\mid{\mathbf{x}}_{t+1}}$ are injective for a bounded function space;
  • B3 (Weakly Monotonic) for any $\dot{\mathbf{z}}_{t},\ddot{\mathbf{z}}_{t}\in\mathcal{Z}^{c}\times\mathcal{Z}^{s}_{t}$ $(\dot{\mathbf{z}}_{t}\neq\ddot{\mathbf{z}}_{t})$, the set $\{\mathbf{x}_{t}:p(\mathbf{x}_{t}|\dot{\mathbf{z}}_{t})\neq p(\mathbf{x}_{t}|\ddot{\mathbf{z}}_{t})\}$ has positive probability, and conditional densities are bounded and continuous;
are satisfied, then ${\mathbf{z}}_{t}$ is block-wise identifiable with regard to $\hat{{\mathbf{z}}}_{t}$ from learned model $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$ under Observation Equivalence.
Illustration of assumptions. The assumptions above are commonly used in the literature on the identification of latent variables under measurement error (hu2008instrumental, ). Firstly, Assumption B1 requires a continuous distribution. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as (yao2022temporally, ).
Overall, the three assumptions impose mild requirements on the underlying data generation process. When these assumptions are satisfied, the block-wise identifiability result established in our theorem holds. Importantly, real-world video scenarios naturally conform to these assumptions, and the structure of many existing generative models is already aligned with the conditions specified in the theorem. A more detailed discussion of these assumptions and their practical relevance can be found in Appendix A.1.
Proof sketch. We separately prove the identifiability of all latent variables and ${\mathbf{z}}^{c}$. For the first part, it is built on fu2025identification , following the line of work from hu2008instrumental . Intuitively, it demonstrates that a minimum of 3 different observations of latent variables are required for identification under the given data generation process. For the second part, we use the contradiction to show that the same ${\mathbf{z}}^{c}$ in different frames of a video can be identified leveraging the invariance.
3.2 Component-wise identification
Theorem 3.3 (Component-wise Identifiability).
Consider video observation $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{T}\}$ generated by process $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$ with latent variables denoted as ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{c}$, according to Equation 1, where ${\mathbf{x}}_{t}\in\mathbb{R}^{n_{x}},{\mathbf{z}}_{t}^{s}\in\mathbb{R}^{n_{s}},{\mathbf{z}}^{c}\in\mathbb{R}^{n_{c}}$. Suppose assumptions in Theorem 3.2 hold. If assumptions
  • C1 (Smooth and Positive Density) the probability density function of latent variables is always third-order differentiable and positive;
  • C2 (Sufficient Changes) let $\eta_{t,i}\triangleq\log p(z^{s}_{t,i}|{\mathbf{z}}_{t-1}^{s})$ and
    $\mathbf{v}_{t,l}\triangleq\Big{(}\frac{\partial^{2}\eta_{t,1}}{\partial z_{t,1}\partial z_{t-1,l}},\cdots,\frac{\partial^{2}\eta_{t,n_{s}}}{\partial z_{t,n}\partial z_{t-1,l}}\Big{)}\oplus\Big{(}\frac{\partial^{3}\eta_{t,1}}{\partial^{2}z_{t,1}\partial z_{t-1,l}},\cdots,\frac{\partial^{3}\eta_{t,n_{s}}}{\partial^{2}z_{t,n}\partial z_{t-1,l}}\Big{)},$ (4)
    for $l\in\{1,2,\cdots,n\}$. For each value of ${\mathbf{z}}_{t}$, there exists $2n_{s}$ different values of $z_{t-1,l}$ such that the $2n_{s}$ vectors ${\mathbf{v}}_{t,l}\in\mathbb{R}^{2n_{s}}$ are linearly independent;
  • C3 (Conditional Independence) the learned $\hat{{\mathbf{z}}}_{t}^{s}$ is independent with $\hat{{\mathbf{z}}}^{c}$, and all entries of $\hat{{\mathbf{z}}}_{t}^{s}$ are mutually independent conditioned on $\hat{{\mathbf{z}}}_{t-1}^{s}$;
are satisfied, then ${\mathbf{z}}_{t}^{s}$ is component-wise identifiable with regard to $\hat{{\mathbf{z}}}_{t}^{s}$ from learned model $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$ under Observation Equivalence.
Proof sketch. In summary, component-wise identification relies on the changeability of style dynamics, i.e., sufficient changes. Starting from the results of block-wise identification, we establish the connection between ${\mathbf{z}}_{t}^{s}$ and $\hat{{\mathbf{z}}}_{t}^{s}$ in terms of their distributions, i.e., $p({\mathbf{z}}_{t})=p(\hat{{\mathbf{z}}}_{t})\cdot|H_{t}|$, where $H_{t}$ is the jacobian matrix. Leveraging the second-order derivative of the log probability, we construct a system of equations with terms of $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}$ and coefficients as specified in assumption C2. We leverage the third-order derivative of the previous latent variable $z_{t-1,l}$ to eliminate $|H_{t}|$, utilizing the fact that the history does not influence the current mapping from estimation to truth. Solving this system yields $\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,j}^{s}}\cdot\frac{\partial z_{t,i}^{s}}{\partial\hat{z}_{t,k}^{s}}=0$.
Illustration of assumptions. The assumptions C1 and C2 on the data generating process are commonly adopted in existing identifiable results for Temporally Causal Representation Learning yao2022temporally . Specifically, C1 implies that the latent variables evolve continuously over time, while C2 ensures that the variability in the data can be effectively captured. The assumption C3 constraints to the learned model requires it to separate different variables of style dynamics into independent parts.
4 Approach
Figure 3: The workflow of the Generator proceeds from left to right, starting with random noise. It first passes through a Temporal Transition Module to obtain a disentangled representation, which is then fed into the synthesis network for frame-level generation. The blue arrow in the figure represents a component-wise flow.
Given our results on identifiability, we implement our model, CoVoGAN. Our architecture is based on StyleGAN2-ADA (Karras2020ada, ), incorporating a Temporal Transition Module in the Generator to enforce the minimal change principle and sufficient change property. Additionally, we add a Video Discriminator to ensure observational equivalence of the joint distribution $p(V)$.
4.1 Model structure
Noise sampling. The structure of the generator is shown in Figure 3. To generate a video with length $T$, we first independently sample random noise from a normal distribution $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We then naively split it into several parts, i.e., $\epsilon=[\epsilon^{c};\epsilon^{s}_{1};\epsilon^{s}_{2};\cdots;\epsilon^{s}_{T}]$.
Temporal Transition Module. Following the generating process in Equation 1, we handle ${\mathbf{z}}^{c}$ and ${\mathbf{z}}_{t}^{s}$ separately. On the one hand, we employ an autoregressive model to capture historical information, followed by a conditional flow to generate ${\mathbf{z}}_{t}^{s}$. Specifically, we implement a Gated Recurrent Unit (GRU) (chung2014empirical, ) and Deep Sigmoid Flow (DSF) (huang2018neural, ), formulated as
${\mathbf{h}}_{t}=\text{GRU}({\mathbf{h}}_{t-1},\epsilon_{t}^{s}),z_{t,i}^{s}=\text{DSF}_{i}(\epsilon_{t,i}^{s};{\mathbf{h}}_{t-1}).$ (5)
On the other hand, since ${\mathbf{z}}^{c}$ are not required to be mutually independent, we use an MLP to generate ${\mathbf{z}}^{c}$, i.e.,
${\mathbf{z}}^{c}=\text{MLP}(\epsilon^{c}).$ (6)
Concatenate ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}^{c}$, and then we obtain the disentangled representation ${\mathbf{z}}_{t}={\mathbf{z}}^{c}\oplus{\mathbf{z}}_{t}^{s}$ for each frame at time step $t$ of the video.
Synthesis network. The synthesis network is designed in the same way as StyleGAN2-ADA. The generated representation ${\mathbf{z}}_{t}$ is first fed into the mapping network to obtain a semantic vector $w({\mathbf{z}}_{t})\in\mathcal{W}$, and then the $t$-th frame of the video is generated by the convolutional network with $w$.
Discriminator structure. To ensure observational equivalence, we implement a video discriminator $D_{V}$ separate from the image discriminator $D_{I}$. For the image discriminator, we follow the design of the original StyleGAN2-ADA. For the video discriminator, we adopt a channel-wise concatenation of activations at different resolutions to model and manage the spatiotemporal output of the generator.
Loss. In addition to the original loss function of StyleGAN2-ADA, we introduce two additional losses: (1) a video discriminator loss, and (2) a mutual information maximization term chen2016infogan between the latent dynamic variables ${\mathbf{z}}_{t}^{s}$ and the intermediate layer outputs of the video discriminator. This encourages the model to learn a more informative and structured representation.
4.2 Relationship between model and theorem.
In this subsection, we discuss the relationship between the CoVoGAN  model and the Theorem.
Block-wise identification. As discussed in Theorem 3.2, achieving block-wise identifiability benefits from minimizing the dimension $n_{s}$ of the style dynamics, especially when the true $n_{s}$ is unknown. In practice, this translates to a hyperparameter selection question. In our experiments, we opted for a relatively modest value of $n_{s}$ and observed that it suffices to attain a satisfactory level of disentanglement capability. Furthermore, as required by the assumptions, the learned variables $\hat{{\mathbf{z}}}_{t}^{s}$ and $\hat{{\mathbf{z}}}^{c}$ are block-wise independent, i.e., $\hat{{\mathbf{z}}}_{t}^{s}\perp\!\!\!\perp\hat{{\mathbf{z}}}^{c}$. This independence is necessary to achieve block-wise identifiability.
Component-wise identification. As outlined in Theorem 3.3, the sufficient change property is a critical assumption for achieving identifiability. To enforce temporally conditional independence, we employ a component-wise flow model that transforms a set of independent noise variables $\epsilon_{t}^{s}$ into the style dynamics, conditioned on historical information. Furthermore, the flow model is designed to maximally preserve the information from $\epsilon_{t,i}^{s}$ to $z_{t,i}^{s}$, enabling the model to effectively capture sufficient variability in the data.
Note that when computing the historical information $h_{t}$, we utilize $\epsilon_{t}^{s}$ instead of ${\mathbf{z}}_{t}^{s}$ (as illustrated in the generating process in Equation 1) as the condition for the component-wise flow. This approach offers two key advantages. First, it simplifies the model architecture since the flow does not need to incorporate the output from another flow. Second, the noise terms already fully characterize the corresponding style dynamics, which remains consistent with the theoretical framework.
Furthermore, given that the precise time lag of dynamic variables remains unspecified a priori in the dataset, the GRU’s gating mechanism can selectively filter out irrelevant historical information that lies outside $\text{Pa}({\mathbf{z}}_{t})$. This capability enables the model to demonstrate significantly superior performance compared to traditional non-gated architectures, such as vanilla RNNs.
A detailed ablation study is presented in Section 5.4.
5 Experiments
In this section, we perform extensive experimental validation to comprehensively evaluate CoVoGAN. We design both qualitative and quantitative experiments to systematically assess the generation quality and controllability of our proposed model. Finally, we perform comprehensive ablation studies to demonstrate the effectiveness of our proposed modules and analyze their individual contributions.
5.1 Experimental setup
Datasets. We evaluate our model on three different real-world datasets: FaceForensics rossler2018faceforensics , SkyTimelapse xiong2018learning , and RealEstate zhou2018stereo . These datasets respectively capture dynamic variations in facial expressions, sky transitions, and complex camera movements, providing comprehensive evaluation scenarios for our method. FaceForensics and SkyTimelapse are widely adopted benchmarks in video generation research, particularly for facial manipulation and natural scene modeling. The RealEstate dataset contains video sequences featuring complex camera movements within static scenes. The camera motions in RealEstate are more intuitive and easily interpretable by human observers, including forward/backward movements, lateral shifts, etc. We particularly leverage this dataset to effectively demonstrate the precise controllability of CoVoGAN in handling explicit camera motion parameters. All datasets consist of videos with a resolution of 256 × 256 pixels, and we employ standard train-test splits for fair evaluation. Detailed statistics and additional information about the datasets are provided in Appendix C.
Evaluation metrics. To comprehensively evaluate the performance of CoVoGAN, we employ both quantitative and qualitative assessment metrics. For quantitative evaluation, we adopt the Fréchet Video Distance (FVD) unterthiner2018towards , a widely-used metric for assessing video generation quality. FVD computes the Wasserstein-2 distance between feature distributions of real and generated videos in a learned feature space, providing a comprehensive measure of visual quality and temporal coherence. We report FVD scores at two different temporal scales: $\text{FVD}_{8}$ and $\text{FVD}_{16}$, where the subscript denotes the number of frames in a video. To better assess the disentanglement capability of our method, we compare SAP kumar2017variational and modularity ridgeway2018learning . Additionally, we measure the Mean Correlation Coefficient (MCC), a standard metric in the Independent Component Analysis (ICA) literature for continuous variables. For qualitative evaluation, we present randomly sampled video sequences generated by our model.
5.2 Video quality
We conduct comparisons between our proposed CoVoGAN and four GAN-based models: MoCoGAN-HD, DIGAN, StyleGAN-V, and MoStGAN-V, across three benchmark datasets: FaceForensics, SkyTimelapse, and RealEstate. We also compare three diffusion-based models: VDM, LVDM he2022latent and Latte ma2024latte . For MoCoGAN-HD, we utilize the image generator pretrained by the original authors and train the remaining components (including the motion generator and discriminator). Due to the unavailability of pretrained models for the RealEstate dataset, we exclude MoCoGAN-HD from the comparisons on this specific dataset. For Latte, we directly use the released checkpoints. For the remaining baselines, we train the models from scratch using their official implementations to ensure comparability. The quantitative results in Table 1 show that CoVoGAN consistently achieves top performance across all datasets, despite diffusion-based models occasionally exhibiting higher visual fidelity. Additional randomly sampled generations are provided in the Appendix E.2.
Figure 4: Disentangled control over identity and motion in (a) FaceForensics, (b) RealEstate.
Table 1: $\text{FVD}_{8}\downarrow$ and $\text{FVD}_{16}\downarrow$ scores across the FaceForensics, SkyTimelapse, and RealEstate datasets.
Method FaceForensics $\text{256}^{2}$ SkyTimelapse $\text{256}^{2}$ RealEstate $\text{256}^{2}$
$\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$
MoCoGAN-HD 140.05 185.51 1214.13 1721.89 - -
DIGAN 57.52 61.65 60.54 105.03 182.86 178.27
StyleGAN-V 49.24 52.70 45.30 62.55 199.66 201.95
MoStGAN-V 47.67 49.85 40.97 55.36 247.77 265.54
VDM 1038.29 1046.60 1099.21 1104.80 1524.17 1526.04
LVDM 136.60 153.38 307.22 319.67 423.54 448.31
Latte 45.49 49.02 40.21 41.84 - -
CoVoGAN (ours) 43.75 48.80 35.58 46.51 154.88 174.87
5.3 Controllability
Block-wise disentanglement. Figure 4 demonstrates the video controllability of CoVoGAN in comparison with baseline methods, highlighting our model’s superior capability for disentanglement between motion and content. The analysis follows a systematic procedure: we first generate a base video sequence, then apply a controlled modification by adding a value to specific motion-related latent variables.
For CoVoGAN, we modify one dimension of the style dynamics ${\mathbf{z}}_{t}^{s}$. For StyleGAN-V and MoStGAN-V, we manipulate one dimension of their (latent) motion code. We apply equivalent modifications to the corresponding latent dimensions in each baseline model. To validate the consistency of our controllability analysis, we randomly sample three distinct video sequences and apply identical
Figure 5: Controllability visualization results on the FaceForensics dataset. Two different concepts of motion are modified to show the component-wise level disentanglement. The first line is a randomly sampled video, and the following lines apply control of eye blinking and head shaking. Each subplot displays a small green arrow indicating the measured distance between the facial contour and the image boundary.
modifications to their respective latent representations. Other baselines are not compared since there are no specific variables with semantic information.
The results show that our proposed CoVoGAN model learns a disentangled representation that effectively separates style dynamics from content elements. (1) This disentanglement enables independent manipulation of motion characteristics while preserving content consistency. (2) A key advantage of this approach is that identical modifications to the style latent space consistently produce similar motion patterns across different content identities.
Baseline models achieve only partial disentanglement, exhibiting two major limitations: (1) visual distortions of the modified videos and (2) inconsistent or misaligned motion patterns when applied to different identities.
Component-wise disentanglement. We further illustrate the precise control over individual motion components. This capability is enabled by the component-wise identifiability of our model, which ensures each latent dimension corresponds to a specific and interpretable motion attribute, e.g., eye blinking or head shaking.
Our experimental procedure begins with randomly sampling two distinct video sequences, as illustrated in the first row of Figure 5. We then selectively modify the latent dimension corresponding to eye blinking dynamics in the second line. Subsequently, we modify a second latent dimension controlling head shaking motion while maintaining the previously adjusted eye blinking pattern in the last line. The results show naturalistic head movements from left to right, synchronized with the preserved eye blinking, illustrating our model’s capability for independent yet coordinated control of multiple motion components.
Table 2: Comparison of disentanglement metrics across different models. Higher is better.
Metrics CoVoGAN StyleGAN-V MoStGAN-V
MCC (%) $\uparrow$ 33.78 29.00 27.95
SAP (%) $\uparrow$ 8.48 4.25 5.90
Modularity (%) $\uparrow$ 17.37 7.66 13.48
Disentanglement metrics. To evaluate disentanglement performance, we first extract semantic annotations from the videos. For this purpose, we conduct experiments on the FaceForensics dataset, which facilitates the extraction of meaningful facial attributes. We utilize a pretrained model from Dlib to extract facial landmarks and compute semantic annotations, including eye size, mouth size, head position, and head angle. We only compare our method with two models that explicitly incorporate semantic representation layers, which we use to compute the metrics. Diffusion-based models without semantic representations are not included in this comparison. For our method, we compute the metrics using the dynamic latent variables $z^{s}_{t}$. As shown in Table 2, our method achieves the best performance across all disentanglement metrics.
5.4 Ablation study
We conduct an ablation study to evaluate the contributions of the proposed components within the Temporal Transition Module, as shown in Table 3.
Replacing the GRU with a standard RNN leads to a noticeable performance drop. This is primarily due to the loss of sparsity provided by the GRU’s gating mechanism, which helps isolate time-delayed effects. Without this mechanism, the model faces a larger search space, making it harder to learn an effective transition function.
Substituting the component-wise flow with a fully connected MLP results in an even more significant degradation. This decline can be attributed to two factors: (1) the mutual independence between style dynamics can no longer be maintained, and (2) capturing sufficient changes becomes more difficult.
Table 3: Ablation study results for different configurations of GRU and component-wise flow on the FaceForensics dataset.
Metric CoVoGAN w/o GRU w/o flow
FVD16 $\downarrow$ 48.80 53.68 82.81
MCC (%) $\uparrow$ 33.78 26.59 8.22
SAP (%) $\uparrow$ 8.48 7.25 0.55
Modularity (%) $\uparrow$ 17.37 12.40 10.24
6 Conclusion
In this paper, we proposed a Temporal Transition Module and implemented it in a GAN to achieve CoVoGAN. By leveraging the principles of minimal and sufficient changes, we successfully disentangled (1) the motion and content of a video, and (2) different concepts within the motion. We established an identifiability guarantee for both block-wise and component-wise disentanglement. Our proposed CoVoGAN model demonstrates high generative quality and controllability. We validated the performance on various datasets and conducted ablation experiments to further confirm the effectiveness of our model. Overall, our work provides a principled and practical framework for disentangled video generation, and opens new directions for fine-grained, interpretable, and controllable visual synthesis. Nonetheless, our study has certain limitations. It primarily centers on theoretical contributions and their empirical validation, while the integration with higher-fidelity generative architectures and state-of-the-art text-to-video synthesis frameworks remains an open direction for future research.
Appendix for
"Controllable Video Generation with Provable Disentanglement"
Appendix A Identifiability Theory
A.1 Proof
Without loss of generality, we first consider the case where $\text{Pa}(z_{t,i}^{s})={\mathbf{z}}_{t-1}^{s}$, meaning that the time-dependent effects are governed by the dynamics of the previous time step.
Theorem A1 (Blockwise Identifiability).
Consider video observation $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{T}\}$ generated by process $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$ with latent variables denoted as ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{c}$, according to Equation 1, where ${\mathbf{x}}_{t}\in\mathbb{R}^{n_{x}},{\mathbf{z}}_{t}^{s}\in\mathbb{R}^{n_{s}},{\mathbf{z}}^{c}\in\mathbb{R}^{n_{c}}$. If assumptions
  • B1 (Positive Density) the probability density function of latent variables is always positive and bounded;
  • B2 (Minimal Changes) the linear operators $L_{{\mathbf{x}}_{t+1}\mid{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}}$ and $L_{{\mathbf{x}}_{t-1}\mid{\mathbf{x}}_{t+1}}$ are injective for bounded function space;
  • B3 (Weakly Monotonic) for any $\dot{\mathbf{z}}_{t},\ddot{\mathbf{z}}_{t}\in\mathcal{Z}^{c}\times\mathcal{Z}^{s}_{t}$ $(\dot{\mathbf{z}}_{t}\neq\ddot{\mathbf{z}}_{t})$, the set $\{\mathbf{x}_{t}:p(\mathbf{x}_{t}|\dot{\mathbf{z}}_{t})\neq p(\mathbf{x}_{t}|\ddot{\mathbf{z}}_{t})\}$ has positive probability, and conditional densities are bounded and continuous;
are satisfied, then ${\mathbf{z}}_{t}$ is blockwisely identifiable with regard to $\hat{{\mathbf{z}}}_{t}$ from learned model $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$ under Observation Equivalence.
Proof.
We first prove the monoblock identification of ${\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}=h(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$, then we prove the blockwise identification ${\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}^{c})$.
Monoblock Identification. Following [10, 9], when assumptions B1, B2, B3 satisfied, the blockwise identifiability of $[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]$ is assured, according to Theorem 3.2(Monoblock identifiability) in [6]. In short ,there exists a invertible function $g$ such that $[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation.
Identification of ${\mathbf{z}}^{c}$. We prove this by contradiction. Suppose that for any $\hat{{\mathbf{z}}}^{c}$, we have
${\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}),$ (7)
where there exist at least two distinct values of $\hat{{\mathbf{z}}}_{t}^{s}$ such that ${\mathbf{z}}^{c}$ takes different values.
When observational equivalence holds, the function remains the same for all values of $t$:
$h_{c}(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})=g_{c}^{-1}({\mathbf{x}}_{t})=(g_{c}^{-1}\circ\hat{g})(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}),$ (8)
where $g_{c}^{-1}$ first demix ${\mathbf{x}}_{t}$ then extract the content part, as defined in Equation 2.
For any two distinct $t\neq t^{\prime}$, we have
$h_{c}(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})={\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}_{t^{\prime}}^{s},\hat{{\mathbf{z}}}^{c}),$ (9)
which holds for all pairs $(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}_{t^{\prime}}^{s})$ within the domain of definition.
According to Assumption B1, the joint distribution $p(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}_{t^{\prime}}^{s})$ is always positive. Thus, to satisfy Equation 9, $\hat{{\mathbf{z}}}_{t}^{s}$ must not contribute to ${\mathbf{z}}^{c}$ through $h_{c}$. In other words, we obtain
${\mathbf{z}}^{c}=h_{c}(\hat{{\mathbf{z}}}^{c}).$ (10)
Theorem A2 (Component-wise Identifiability).
Consider video observation $V=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{T}\}$ generated by process $(g,\textbf{f}^{s},\textbf{f}^{c},\textbf{p}^{s},\textbf{p}^{c})$ with latent variables denoted as ${\mathbf{z}}_{t}^{s}$ and ${\mathbf{z}}_{c}$, according to Equation 1, where ${\mathbf{x}}_{t}\in\mathbb{R}^{n_{x}},{\mathbf{z}}_{t}^{s}\in\mathbb{R}^{n_{s}},{\mathbf{z}}^{c}\in\mathbb{R}^{n_{c}}$. Suppose assumptions in Theorem A1 hold. If assumptions
  • C1 (Smooth and Positive Density) the probability density function of latent variables is always third-order differentiable and positive;
  • C2 (Sufficient Changes) let $\eta_{t,i}\triangleq\log p(z^{s}_{t,i}|{\mathbf{z}}_{t-1}^{s})$ and
    $\mathbf{v}_{t,l}\triangleq\Big{(}\frac{\partial^{2}\eta_{t,1}}{\partial z_{t,1}\partial z_{t-1,l}},\cdots,\frac{\partial^{2}\eta_{t,n_{s}}}{\partial z_{t,n}\partial z_{t-1,l}}\Big{)}\oplus\Big{(}\frac{\partial^{3}\eta_{t,1}}{\partial^{2}z_{t,1}\partial z_{t-1,l}},\cdots,\frac{\partial^{3}\eta_{t,n_{s}}}{\partial^{2}z_{t,n}\partial z_{t-1,l}}\Big{)},$ (11)
    for $l\in\{1,2,\cdots,n\}$. For each value of ${\mathbf{z}}_{t}$, there exists $2n_{s}$ different of values of $z_{t-1,l}$ such that the $2n_{s}$ vector ${\mathbf{v}}_{t,l}\in\mathbb{R}^{2n_{s}}$ are linearly independent;
  • C3 (Conditional Independence) the learned $\hat{{\mathbf{z}}}_{t}^{s}$ is independent with $\hat{{\mathbf{z}}}^{c}$, and all entries of $\hat{{\mathbf{z}}}_{t}^{s}$ are mutually independent conditioned on $\hat{{\mathbf{z}}}_{t-1}^{s}$;
are satisfied, then ${\mathbf{z}}_{t}^{s}$ is component-wisely identifiable with regard to $\hat{{\mathbf{z}}}_{t}^{s}$ from learned model $(\hat{g},\hat{\textbf{f}}^{s},\hat{\textbf{f}}^{c},\hat{\textbf{p}}^{s},\hat{\textbf{p}}^{c})$ under Observation Equivalence.
Proof.
According to Theorem A1, we have
$[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{\mathbf{z}}_{t}^{s},\hat{\mathbf{z}}^{c}]),$ (12)
where [.][.][ . ] denotes the concatenation operation. The corresponding Jacobian matrix can be formulated as
$H_{t}=\begin{bmatrix}\frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}_{t}^{s}}&\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}_{t}^{s}}\\ \frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}^{c}}&\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}^{c}}\end{bmatrix}.$ (13)
Consider a mapping from $({\mathbf{x}}_{t-1},\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$ to $({\mathbf{x}}_{t-1},{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c})$ and its Jacobian matrix
$\begin{bmatrix}\mathbf{I}&\mathbf{0}\\ \mathbf{*}&H_{t}\end{bmatrix},$ (14)
where * stands for any matrix, and the absolute value of the determination of this Jacobian is $|H_{t}|$. Therefore $p({\mathbf{x}}_{t-1},{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c})=p({\mathbf{x}}_{t-1},\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})/|H_{t}|$. Dividing both side by $p({\mathbf{x}}_{t-1})$ gives
[…]
For the first two implications, we utilize the inversion of the mixing function to replace the condition. For the third, since ${\mathbf{z}}^{c}$ is conditioned on itself, it remains fixed. For the last one, given that in the generating process, $\epsilon_{t,i}^{s}$ and $\epsilon_{j}^{c}$ for all $i,j,t$ are independently sampled, their disjoint successors are also independent, i.e., ${\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s}\perp{\mathbf{z}}^{c}$. Similarly, following assumption C3, we have $\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s}\perp\hat{{\mathbf{z}}}^{c}$. Thus, we can remove ${\mathbf{z}}^{c}$ from the condition.
For simplicity, denote $\mathbb{\eta}_{t}\triangleq\log p({\mathbf{z}}^{s}_{t}|{\mathbf{z}}_{t-1}^{s})$ and $\eta_{t,i}\triangleq\log p(z^{s}_{t,i}|{\mathbf{z}}_{t-1}^{s})$ and we have
$\mathbb{\eta}_{t}=\hat{\mathbb{\eta}}_{t}-\log|H_{t}|.$ (16)
For any two different $\hat{z}_{t,i}^{s},\hat{z}_{t,j}^{s}\in{\mathbf{z}}_{t}^{s}$, in partial derivative with regard to $\hat{z}_{t,i}^{s}$ gives
$\sum_{k=1}^{n_{s}}\frac{\partial\mathbb{\eta}_{t}}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}=\frac{\partial\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}}-\frac{\partial\log|H_{t}|}{\partial\hat{z}_{t,i}^{s}}.$ (17)
Reorganize the left-hand side of Equation 17 with mutual independence of ${\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s}$ yields
$\sum_{k=1}^{n_{s}}\frac{\partial\mathbb{\eta}_{t}}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}=\sum_{k=1}^{n_{s}}\frac{\partial\prod_{k^{\prime}=1}^{n_{s}}\eta_{t,k^{\prime}}}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}=\sum_{k=1}^{n_{s}}\frac{\partial\eta_{t,k}}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}},$ (18)
and we have
$\sum_{k=1}^{n_{s}}\frac{\partial\eta_{t,k}}{\partial z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}=\frac{\partial\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}}-\frac{\partial\log|H_{t}|}{\partial\hat{z}_{t,i}^{s}}.$ (19)
Further get the second-order derivative with regard to $\hat{z}_{t,j}^{s}$ as
$\sum_{k=1}^{n_{s}}\frac{\partial^{2}\eta_{t,k}}{\partial^{2}z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}^{s}}+\sum_{k=1}^{n_{s}}\frac{\partial\eta_{t,k}}{\partial z_{t,k}^{s}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=\frac{\partial^{2}\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}-\frac{\partial^{2}\log|H_{t}|}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}.$ (20)
Next, using the mutual independence of $\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s}$ in assumption C3, we have $\frac{\partial^{2}\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=0$ according to the connection between conditional independence and cross derivatives [24]. Thus we have
$\sum_{k=1}^{n_{s}}\frac{\partial^{2}\eta_{t,k}}{\partial^{2}z_{t,k}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}^{s}}+\sum_{k=1}^{n_{s}}\frac{\partial\eta_{t,k}}{\partial z_{t,k}^{s}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=-\frac{\partial^{2}\log|H_{t}|}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}.$ (21)
Now we get the third-order derivative with regard to any $z_{t-1,l}^{s}$ as
$\sum_{k=1}^{n_{s}}\frac{\partial^{3}\eta_{t,k}}{\partial^{2}z_{t,k}^{s}\partial z_{t-1,l}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}^{s}}+\sum_{k=1}^{n_{s}}\frac{\partial\eta_{t,k}}{\partial z_{t,k}^{s}}\cdot\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=0,$ (22)
where we use the property that the entries of $H_{t}$ do not depend on $z_{t-1,l}^{s}$.
Given assumption C2, there exists $2n_{s}$ different values of $z_{t-1,l}^{s}$ such that the $2n_{s}$ vectors ${\mathbf{v}}_{t,l}$ linearly independent. The only solution to Equation 22 is to set
$\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}}\cdot\frac{\partial z_{t,k}^{s}}{\partial\hat{z}_{t,j}^{s}}=0,\frac{\partial^{2}z_{t,k}^{s}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=0.$ (23)
According to Theorem A1, the blockwise identifiability is established. Thus,
$H_{t}=\begin{bmatrix}\frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}_{t}^{s}}&\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}_{t}^{s}}\\ \frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}^{c}}&\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}^{c}}\end{bmatrix}$ (24)
is invertible, with $\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}_{t}^{s}}=0$ and $\frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}_{t}^{s}}$ has at most one nonzero element in each row and each column.
Thus, we have
$H_{t}=\begin{bmatrix}\frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}_{t}^{s}}&0\\ \frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}^{c}}&\frac{\partial{\mathbf{z}}^{c}}{\partial\hat{\mathbf{z}}^{c}}\end{bmatrix}$ (25)
and $\frac{\partial{\mathbf{z}}_{t}^{s}}{\partial\hat{\mathbf{z}}_{t}^{s}}$ must have one and only one non-zero entry in each column and row.
A.2 Discussion of Assumptions
In this section, we give a brief discussion of the assumptions in the real world scenarios.
Assumption B1 requires a smooth distribution, which is usually hold in the real world. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as [52]. The aforementioned assumptions concern the underlying data-generating process. The aforementioned assumptions pertain to the underlying data-generating process, and are often easily satisfied in real-world video scenarios. Given a model that can explicitly disentangle motion and identity while maintaining high generative fidelity, block-wise disentanglement becomes a natural and attainable property under these conditions. This also explains why models such as MoCoGAN [37] are able to effectively separate motion and identity.
Assumption C1 further requires the underlying distribution to be smooth, while Assumptions C2 and C3 impose mild conditions on the variability of latent factors. These assumptions are not overly restrictive and are often satisfied in practice. Notably, even when the sufficiency conditions are satisfied only by a subset of latent variables, or when mutual independence is partially violated, identifiability can still be achieved at the subspace level, as demonstrated by [17] in Section A.1. To further attain disentanglement at the component-wise level, the architectural design of our proposed Temporal Transition Module (TTM) is essential.
Appendix B Related Works
B.1 Controllable Video Generation
Recent advances in controllable video generation have led to significant progress, with text-to-video (T2V) models [51, 34, 8, 57, 56] achieving impressive results in generating videos from textual descriptions. However, effectiveness of the control is highly dependent on the quality of the input prompt, making it difficult to achieve fine-grained control over the generated content. An alternative method for control involves leveraging side information such as pose [36, 59], camera motion [50, 45], depth [27, 48] and so on. While this approach allows for more precise control, it requires paired data, which can be challenging to collect. Besides, most of the aforementioned alignment-based techniques share a common issue: the control signals are directly aligned with the entire video. This issue not only reduces efficiency but also complicates the task of achieving independent control over different aspects of the video, which further motivates us to propose a framework to find the disentanglement representation for conditional generation.
B.2 Nonlinear Independent Component Analysis
Nonlinear independent component analysis offers a potential approach to uncover latent causal variables in time series data. These methods typically utilize auxiliary information, such as class labels or domain-specific indices, and impose independence constraints to enhance the identifiability of latent variables. Time-contrastive learning (TCL) [12] builds on the assumption of independent sources and takes advantage of the variability in variance across different segments of data. Similar Permutation-based contrastive learning (PCL) [13] introduces a learning framework that tell true independent sources from their permuted counterparts. Additionally, i-VAE [16] employs deep neural networks and Variational Autoencoders (VAEs) to closely approximate the joint distribution of observed data and auxiliary non-stationary regimes. Recently, [53, 52] extends the identifiability to linear and nonlinear non-Gaussian cases without auxiliary variables, respectively. CaRiNG [2] further tackles the case when the mixing process is non-invertible. Additionally, CITRIS [26, 25] emphasizes the use of intervention target data, and IDOL [22] incorporates sparsity into the latent transition process to identify latent variables, even in the presence of instantaneous effects.
Appendix C Datasets Details
FaceForensics [32] is a forensics dataset consisting of 1000 original video sequences all videos contain a trackable mostly frontal face without occlusions. SkyTimelapse [49] typically consists of sequential images or videos capturing the dynamic behavior of the sky over time.
RealEstate10K [58] is a large dataset of camera poses corresponding to 10 million frames derived from about 80,000 video clips, gathered from about 10,000 YouTube videos. For each clip, the poses form a trajectory where each pose specifies the camera position and orientation along the trajectory. These poses are derived by running SLAM and bundle adjustment algorithms on a large set of videos. To the best of our knowledge, the proposed method is the first GAN-based approach to leverage the complex camera pose dataset for validating unconditional video generation.
Appendix D Reproductability
All of our models are trained on an NVIDIA A100 40G GPU. Our model requires around a throughput of 20 million frames for convergence. For the baseline models, we use the official implementation with the default hyperparameters. The configuration of hyperparameters for our model training is as shown in Table  4.
Table 4: Configuration details for the model and training setup.
Name Variable Name Value/Description
Information regularization term weight lambda_KL 1
Dimensionality of ${\mathbf{z}}^{c}$ 512
Dimensionality of GRU - 64
Dimensionality of ${\mathbf{z}}_{s}^{t}$ FaceForensics: 4, SkyTimelapse: 12, RealEstate: 12
Batch size (per GPU) batch 16
Conditional mode cond_mode flow
Flow normalization flow_norm 1
Sparsity weight lambda_sparse 0.1
Number of input channels channel 3
Number of mapping layers num_layers 8
Label embedding features embed_features 512
Intermediate layer features layer_features 512
Activation function activation lrelu
Learning rate multiplier lr_multiplier 0.01
Moving average decay w_avg_beta 0.995
Discriminator architecture - resnet
Channel base channel_base 32768
Maximum number of channels channel_max 512
The architecture of our proposed Temporal Transition Module is shown in Table 5.
Table 5: Architecture of Temporal Transition Module
Component Structure
MappingNetwork.gru $4\times 256\times 3$
MappingNetwork.h_to_c FullyConnectedLayer ($256\times 64$)
MappingNetwork.embed FullyConnectedLayer ($64\times 512$)
MappingNetwork.flow.model.0 DenseSigmoidFlow
MappingNetwork.flow.model.1 DenseSigmoidFlow
MappingNetwork.flow_fc0 FullyConnectedLayer ($512\times 512$)
MappingNetwork.flow_fc1 FullyConnectedLayer ($512\times 512$)
MappingNetwork.flow_fc2 FullyConnectedLayer ($512\times 284$)
FullyConnectedLayer FullyConnectedLayer ($512\times 512$)
FullyConnectedLayer FullyConnectedLayer ($512\times 512$)
Appendix E More Experiments
E.1 Computational Efficiency.
We also compare the computational efficiency with that of the baselines. The size of our model’s generator is comparable to that of StyleGAN2-ADA but significantly smaller than the generators of the baselines. Additionally, the inference time of our model is much faster, as shown in Table 6.
Table 6: Comparison of different models in terms of generator parameters (Params, in millions) and the time (Time, in seconds) required to generate 16-frame 2048 videos with a resolution of $256\times 256$.
Method Params (M) Time (s)
StyleGAN2-ADA 23.19 -
DIGAN 69.97 142.06
StyleGAN-v 32.11 181.45
MoStGAN 40.92 207.11
Latte 757.38 45340.74
CoVoGAN (ours) 24.98 98.93
E.2 More Videos
In this section, we provide more qualitative video results generated by our approach. As shown in Figure 6, we compare the generative quality of the videos among all models. As can be seen in Figure 7, our method can control different identities of sky scenes with consistent constructed motions.
Figure 6: Random samples from the comparison baselines and our model on real-world datasets FaceForensics $256^{2}$, SkyTimelapse $256^{2}$, RealEstate $256^{2}$. Start from t = 0 and report every second frame from a 16-frame video clip.
Figure 7: Generate same motion of different identities on SkyTimelapse Dataset.
Appendix F Impact Statements
This study introduces both a theoretical framework and a practical approach for extracting disentangled causal representations from videos. Such advancements enable the development of more transparent and interpretative models, enhancing our grasp of causal dynamics in real-world settings. This approach may benefit many real-world applications, including healthcare, auto-driving, content generation, marketing and so on.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: A detailed list of contribution can be found at the end of Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Assumptions mentioned in the theorem part are discussed in the ’Illustration of assumption’ part and Appendix A.2. A discussion of limitation can be found at the end of Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Proof of the theorem can be found in Appendix A.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code is provided in the supplemental material. The details of the implementation can be found in the Appendix D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release the code in the supplemental materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting can be found in the Section 5. The details of the implementation can be found in the Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive for a video generative model.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The required computing resources are provided in the Appendix D and E.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The impact statement can be found in Appendix F.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original paper of baselines and datasets used in this paper are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.